OpenStudy (anonymous):

Identify the following series as arithmetic, geometric, both, or neither. 1/4 + 2/5 + 3/6 + 4/7 I know that the general term is n/n+3 but because the difference isn't something concrete, for example.. 4,6,8,10 d=2, would it be neither?

OpenStudy (shubhamsrg):

nep..neither..

OpenStudy (shubhamsrg):

i meant yep**

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (jamesj):

But do you know why it is neither?

OpenStudy (jamesj):

A series is arithmetic if and only if the difference of each pair of successive terms is equal. You can see the difference 2/5 - 1/4 = 3/20 and 3/6 - 2/5 = 1/10 so the series isn't arithmetic. Similarly, a series is geometric if and of only the ratio of each of pair of successive terms is equal. \[ \frac{2/5}{1/4} = \frac{8}{5}\] but \[ \frac{3/6}{2/5} = \frac{5}{4} \] Hence the series is not geometric either.

OpenStudy (anonymous):

Oh. That is very helpful! Thanks :) And for the geometric, what if it was like a sequence of 1,4,6,8. That wouldn't be geometric either because 1/4 is not equal to 3/4(6/8 reduced)?